OpenStudy (anonymous):

solve by substitution method -x+5y=-6 3x-5y=-2

OpenStudy (anonymous):

is ur question correct

OpenStudy (anonymous):

yes, i got no solution

OpenStudy (anonymous):

ya i got no solution either

jhonyy9 (jhonyy9):

why ??? there are two solutions ... so after assum will get 2x=-8 so x=-4 than 4+5y= -6 so 5y= -10 so y= -2 prove : -x+5y=-6 than -(-4)+5(-2)=4-10=-6 so than we get -6=-6 - right ???